# MT2KD: Towards A General-Purpose Encoder for Speech, Speaker, and Audio Events

### v1

@@ Abstract
With the advances in deep learning, the performance of end-to-end (E2E) single-task models for speech and audio processing has been constantly improving. However, it is still challenging to build a general-purpose model with high performance on multiple tasks, since different speech and audio processing tasks usually require different training data, input features, or model architectures to achieve optimal performance. In this work, MT2KD, a novel two-stage multi-task learning framework is proposed to build a general-purpose speech and audio encoder that jointly performs three fundamental tasks: automatic speech recognition (ASR), audio tagging (AT) and speaker verification (SV). In the first stage, multi-teacher knowledge distillation (KD) is applied to align the feature spaces of three single-task high-performance teacher encoders into a single student encoder using the same unlabelled data. In the second stage, multi-task supervised fine-tuning is carried out by initialising the model from the first stage and training on the separate labelled data of each single task. Experiments demonstrate that the proposed multi-task training pipeline significantly outperforms a baseline model trained with multi-task learning from scratch. The final system achieves good performance on ASR, AT and SV: with less than 4% relative word-error-rate increase on ASR, only 1.9 lower mean averaged precision on AT and 0.23% absolute higher equal error rate on SV compared to the best-performing single-task encoders, using only a 66M total model parameters.
@@ I Introduction
Although high performance on each task can be achieved using separate single-task models, it is still desirable to construct a single model performing multiple speech and audio processing tasks. Such a model could have the following advantages. First, a multi-task model is would be computation-efficient and parameter-efficient since the computation and parameters are shared by multiple tasks. Second, multi-task learning may create task synergy, leading to improved performance over single-task models. Third, with the breakthrough in the development of artificial general intelligence [2], general-purpose audio models are of increasing interest due to their ability to unify multiple tasks [3, 4, 5]. However, multi-task learning of different speech and audio processing tasks faces the following issues. First, the model architecture used is different across different tasks. Task-specific architectures are employed in single-task models to achieve optimal performance and using a single model architecture could lead to sub-optimal performance on each task. Second, speech processing and audio processing tasks require different training data, and coordinating multi-task learning is challenging. Third, the objective of different tasks could interfere with each other, leading to reduced task synergy. For example, an automatic speech recognition (ASR) model should output the same text regardless of the speaker identity, whereas speaker verification (SV) aims at distinguishing different speakers regardless of the text spoken. Recently, self-supervised learning (SSL) leveraging vast amounts of unlabelled speech has achieved promising results on various speech processing tasks using the same pre-trained backbone encoder [4, 6], indicating that the feature space of various speech processing tasks can be aligned. Whisper [7] performs multi-task training of two speech processing tasks: ASR and automatic speech translation. However, other audio processing tasks are not included in its training pipeline.
@@ II-B Audio Tagging
In AT, a model receives an audio clip and predicts the type of sound event. Since a single audio clip can be categorised into multiple different classes, AT is usually treated as a multi-class classification task. Due to the capability of capturing short term information, convolutional neural networks (CNNs) achieved good performance [18] and became a popular backbone model for AT [18, 19, 20]. Gong et al. [21] adopted a transformer model pre-trained on image classification task as the initialisation and trained the model on audio tagging data. The proposed audio spectrogram transformer (AST) adapts quickly to AT and outperforms previous CNN-based networks. AST splits the input log-mel filterbank features into square patches spanning over both time and frequency domain. An embedding is extracted for each patch, which is fed as input to the transformer backbone model. BEATs [22] applied self-supervised pre-training for AT by generating discrete audio tokens via random projection quantisation [23]. The pre-training of the quantizer and the backbone model is performed iteratively using unlabelled data. The backbone model follows the architecture in AST [21]. In [22], the log-Mel filterbank features are split into $N$ 16x16 patches with an overlap of 6 in both time and frequency domains. The patch embeddings are extracted from each patch, which are fed to the transformer encoder as a sequence of token embeddings. The final prediction is made by averaging the logits $\textbf{p}_{1:N}$ of each patch followed by a sigmoid activation. The training loss $\mathcal{L}_{AT}$ for AT is formulated as follows:
@@ II-C Speaker Verification
SV is the task of determining if some speech is from a particular speaker identity, which is usually accomplished by measuring the similarity between the speaker embeddings extracted from the segment with an enrolled voice print. As for ASR, filterbank features are commonly used as input to SV. CNN-based networks demonstrated good performance on SV [24, 25, 26] due to their ability of capturing local patterns. ECAPA-TDNN is one of the most popular architectures, which consists of a frame-level features extractor (encoder) and a segment-level statistical self-attentive pooling module to aggregate the speaker embedding at the segment level. The encoder extracts high-dimensional frame-level acoustic features given the input features and the pooling module adopts a channel-dependent attention mechanism and re-weights the frame-level acoustic features. The final speaker embedding is obtained by projecting the mean and standard deviation aggregated from the weighted frame-level embeddings to a fixed-dimensional vector. During training, a classification layer is apppended after the speaker embedding for predicting the speaker id and the whole model is trained with the cross-entropy loss. When performing speaker verification, the classification layer is discarded and only the speaker embeddings are used. If the cosine similarity between two speaker embeddings exceeds a pre-defined threshold, the two speakers are considered identical.
@@ II-D Multi-task Learning of Speech and Audio Processing
There has been a significant amount of work on multi-task training for speech processing [29]. Multi-task training of speaker recognition and speech recognition is accomplished by utilising the inter-task features extracted from the task-specific branch [30]. Whisper is trained on various speech-processing tasks, including multilingual ASR, speech translation (X $\rightarrow$ English), spoken language identification and voice activity detection. It is demonstrated that multi-task training facilitates the learning of English ASR as the model size and training data scale up.
However, few attempts have been made to build a multi-task model for both audio and speech processing tasks since they require very different training data and different model architectures. Whisper-AT [17] discovered that the audio encoder of the supervised-trained Whisper [7] is highly correlated to non-speech sounds and integrated the AT task in the encoder of Whisper. The layer-wise embeddings of the Whisper encoder are combined with a relatively small network in order to perform audio tagging. When fine-tuned from a Whisper model, the model achieved a mean average precision (mAP) of 45.6 on AudioSet, surpassing existing models using frame-level audio features as input. However, if the encoder parameters are frozen so that the ASR capability is unaffected, the mAP drops to 41.3. Nonetheless, it still indicates that audio and speech processing tasks can be accomplished simultaneously.
@@ II-E Knowledge Distillation in Speech and Audio Processing
KD [31], also known as teacher-student learning, is a widely used technique for model compression and performance improvement. In KD, the student model is trained to match the output of the teacher model, and different KD loss function can be used depending on the form of teacher output, also called as teacher labels. Depending on how and when the teacher labels are generated, KD can be divided into two categories: off-line KD and on-line KD. In off-line KD, the teacher labels are pre-computed and stored at the data preprocessing stage, and are only accessed during KD training. In on-line KD, the teacher labels are computed on-the-fly, requiring extra computation and GPU memory consumption compared to off-line KD since the teacher model need to be forwarded in each batch. However, it also reduces the storage cost, especially for large corpora. Furthermore, consistent data augmentation [32] is easier to implement in on-line KD as long as the teacher and student share the same input data.
@@
In AT, KD has been found to yield better performance than training on the ground-truth multi-hot labels [14]. CED [14] adopts consistent KD in an off-line manner and utilised an ensemble teacher model and achieved state-of-the-art mAP of 50.0 on AudioSet [36]. To reduce the cost of storing augmented input audio features and teacher labels, the random seed for data augmentation and the corresponding teacher top-K logits are stored.
@@ III-A Motivation
    Different tasks can requires different types of input. For example, typical ASR or SV systems require frame-level input features, whereas state-of-the-art AT systems operates on the patch-level input at much higher granularity.
@@ III-C Multi-teacher Multi-task KD Pre-training
where $\mathcal{M}_{\text{ASR}}$ is a module transforming the student encoder representation to the same dimension of the teacher targets. Since the teacher AT model operates on the patch-level features, performing frame-level KD on AT using the teacher embeddings as the target as in ASR is infeasible. Due to the flexibility of KD, the K-dimensional output $C^{\mathcal{T}}$ of the logits before the final Sigmoid activation of $\mathcal{T}^{\text{AT}}$ can be used as teacher labels, where $K$ is the number of audio events. To perform AT, a dedicated module $\mathcal{M}_{\text{AT}}$ that takes the encoder outputs at $i_{\text{AT}}$-th layer also generates a C-dimensional logits. The binary cross-entropy loss between the teacher and student prediction is computed as follows:
@@
where $\mathbf{v}$ is the speaker embeddings predicted by the teacher model and CosineSim is the cosine similarity between two un-normalized vectors.
@@
where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are the scaling factors for the loss functions of ASR, AT and SV, respectively. Despite that there is no dedicated multi-task dataset for ASR, SV and AT, computing three losses for one sample is still theoretically possible because no ground truth labels are not needed during KD. However, only one loss based on the task from which the sample comes can also be computed.
@@ III-D Multi-task Fine-tuning
For audio tagging, the weights of the linear layer projecting the encoder features to number of classes of audio events in the pre-training stage is kept as initialisation. During fine-tuning, the binary cross-entropy loss is adopted as during pre-training. For speaker verification, a linear layer projecting the speaker embeddings to number of speakers is used and the cross-entropy loss is adopted.
@@ IV-A Teacher Models
The encoder of Whisper large-v3 [7], BEATs [22] and ECAPA-TDNN [25] are adopted as the teacher model for ASR, AT and SV, respectively. The encoder of the Whisper large-v3 model [^1] https://github.com/openai/whisper is a transformer model consisting of 24 transformer layers. It receives 128-channel log-magnitude Mel spectrogram representation computed on 25ms windows with a stride of 10ms and generates 1280-D speech representations at a frame-rate of approximately 50 Hz. The Whisper model was trained on over 680,000 hours of weak supervised data to perform ASR and automatic speech translation. Since Whisper is trained on a massive amount of supervised speech data under diverse acoustic conditions, its encoder representations could be beneficial for other tasks. The BEATs models is an audio spectrogram transformer (AST) [21], whose inputs are embeddings extracted from 16x16 patches on the 128-D mel-filter bank features with an overlap of 6. The BEATs checkpoint [^2] https://github.com/microsoft/unilm/tree/master/beats#beats is pre-trained and fine-tuned on AudioSet 2M. The ECAPA-TDNN model adopts a time delay neural network as backbone and has a self-attentive pooling module to extract speaker information. The checkpoint [^3] https://huggingface.co/speechbrain/spkrec-ecapa-voxceleb trained on VoxCeleb1 [37] and VoxCeleb2 [38] was adopted. The model receives 80-D filterbank features as input and produces a single 192-D speaker embedding for a clip of audio. The performance of the teacher models are shown in Table II.
@@ IV-B Student Model
During fine-tuning, the $\mathcal{M}_{\text{AT}}$ is not modified and the binary cross-entropy loss is adopted as the loss for AT training with labelled data. An additional classification layer is added after $\mathcal{M}_{\text{SV}}$ for speaker classification and cross-entropy loss is used for SV training. A predictor and joiner of the neural transducer are added onto the pre-trained student encoder for ASR. Note that the projection layer for learning Whisper embeddings is also kept during fine-tuning. Pruned-RNNT [41] loss, an efficient variant of the RNN-T loss function is used for ASR.
@@ IV-D Training Details
Training is terminated in the first stage of KD pre-training after iterating over the LS-960 or LS-100 for 90 times. In multi-task pre-training and fine-tuning, the LibriSpeech dataset is repeated to match the amount of data from the other two tasks. The dataset configuration is shown in Table I. When AS-2M is used during multi-task training, LS-960 is repeated 5 times. When performing ASR and SV multi-task training, LS-960 is repeated twice. During pre-training, the last ten checkpoints are averaged and used as the initialisation for fine-tuning. After fine-tuning, the results are reported also on the average of the last ten checkpoints.
@@
Two 32GB NVIDIA-V100 GPUs were used in the small scale experiments, while the other experiments used four 32GB NVIDIA-V100 GPUs. Each batch has approximately 1000 seconds of audios. The whole training and fine-tuning process was implemented using icefall [^4] https://github.com/k2-fsa/icefall and Lhotse [42] was used for data preparation.
@@ V-B Single-task KD pre-training
The results of performing single-task KD training of the student model are shown in Table IV. Due to the pre-training loss selection for AT and SV, the student model is already capable of performing AT and SV. Therefore, mAP and EER are used as the evaluation metric for AT and SV. The L1 loss between the teacher Whisper and student embeddings computed on the dev-clean and dev-other set is used to measure the KD training performance for ASR. As shown in Table IV, the KD performance improves as the amount of training data grows. The student model trained with KD using the unlabelled AS-2M outperforms the model trained with the same amount of labelled data, which is consistent with the findings in  [14, 12]. Since AudioSet is a multi-class classification task and many of the classes are correlated, the soft teacher label conveys richer information than the hard label. This also implies that AT models leveraging frame-level features are capable of achieving good results. For SV, the student model trained with KD obtained a lower EER than training from scratch with labelled data. We hypothesise that learning the speaker embedding directly by maximising the cosine similarity between the teacher and student speaker embeddings is an easier and more direct objective than classifying a speaker.

### v3

@@ Abstract
With the advances in deep learning, the performance of end-to-end (E2E) single-task models for speech and audio processing has been constantly improving. However, it is still challenging to build a general-purpose model with high performance on multiple tasks, since different speech and audio processing tasks usually require different training data, input features, or model architectures to achieve optimal performance. In this work, MT2KD, a novel two-stage multi-task learning framework is proposed to build a general-purpose speech and audio encoder that jointly performs three fundamental tasks: automatic speech recognition (ASR), audio tagging (AT) and speaker verification (SV). In the first stage, multi-teacher knowledge distillation (KD) is applied to align the feature spaces of three single-task high-performance teacher encoders into a single student encoder using the same unlabelled data. In the second stage, multi-task supervised fine-tuning is carried out by initialising the model from the first stage and training on the separate labelled data of each single task. Experiments demonstrate that the proposed multi-task training pipeline significantly outperforms a baseline model trained with multi-task learning from scratch. The final system achieves good performance on ASR, AT and SV: with less than 4% relative word-error-rate increase on ASR, only 1.9 lower mean averaged precision on AT and 0.23% absolute higher equal error rate on SV compared to the best-performing single-task encoders, using only a total of 66M model parameters.
@@ I Introduction
Although high performance on each task can be achieved using separate single-task models, it is still desirable to construct a single model performing multiple speech and audio processing tasks. Such a model could have the following advantages. First, a multi-task model is both computation-efficient and parameter-efficient since the computation and parameters are shared by multiple tasks. Second, multi-task learning may create task synergy, leading to improved performance over single-task models. Third, with the breakthrough in the development of artificial general intelligence [2], general-purpose audio models are of increasing interest due to their ability to unify multiple tasks [3, 4, 5]. However, multi-task learning of different speech and audio processing tasks faces the following issues. First, the model architecture used is different across different tasks. Task-specific architectures are employed in single-task models to achieve optimal performance and using a single model architecture could lead to sub-optimal performance on each task. Second, speech processing and audio processing tasks require different training data, and coordinating multi-task learning is challenging. Third, the objective of different tasks could interfere with each other, leading to reduced task synergy. For example, an automatic speech recognition (ASR) model should output the same text regardless of the speaker identity, whereas speaker verification (SV) aims at distinguishing different speakers regardless of the text spoken. Recently, self-supervised learning (SSL) leveraging vast amounts of unlabelled speech has achieved promising results on various speech processing tasks using the same pre-trained backbone encoder [4, 6], indicating that the feature space of various speech processing tasks can be aligned. Whisper [7] performs multi-task training of two speech processing tasks: ASR and automatic speech translation. However, other audio processing tasks are not included in its training pipeline.
@@ II-B Audio Tagging
In AT, a model receives an audio clip and predicts the type of sound event. Since a single audio clip can be categorised into multiple different classes, AT is usually treated as a multi-class classification task. Due to the capability of capturing short term information, convolutional neural networks (CNNs) achieved good performance [18] and became a popular backbone model for AT [18, 19, 20]. Gong et al. [21] adopted a transformer model pre-trained on image classification task as the initialisation and trained the model on audio tagging data. The proposed audio spectrogram transformer (AST) adapts quickly to AT and outperforms previous CNN-based networks. AST splits the input log-mel filterbank features into square patches spanning over both time and frequency domain. An embedding is extracted for each patch, which is fed as input to the transformer backbone model. BEATs [22] applied self-supervised pre-training for AT by generating discrete audio tokens via random projection quantisation [23]. The pre-training of the quantiser and the backbone model is performed iteratively using unlabelled data. The backbone model follows the architecture in AST [21]. In [22], the log-Mel filterbank features are split into $N$ 16x16 patches with an overlap of 6 in both time and frequency domains. The patch embeddings are extracted from each patch, which are fed to the transformer encoder as a sequence of token embeddings. The final prediction is made by averaging the logits $\textbf{p}_{1:N}$ of each patch followed by a sigmoid activation. The training loss $\mathcal{L}_{AT}$ for AT is formulated as follows:
@@ II-C Speaker Verification
SV is the task of determining if some speech is from a particular speaker, which is usually accomplished by measuring the similarity between the speaker embeddings extracted from the segment with an enrolled voice print. As for ASR, filterbank features are commonly used as input to SV. CNN-based networks demonstrated good performance on SV [24, 25, 26] due to their ability of capturing local patterns. ECAPA-TDNN is one of the most popular architectures, which consists of a frame-level features extractor (encoder) and a segment-level statistical self-attentive pooling module to aggregate the speaker embedding at the segment level. The encoder extracts high-dimensional frame-level acoustic features given the input features and the pooling module adopts a channel-dependent attention mechanism and re-weights the frame-level acoustic features. The final speaker embedding is obtained by projecting the mean and standard deviation aggregated from the weighted frame-level embeddings to a fixed-dimensional vector. During training, a classification layer is apppended after the speaker embedding for predicting the speaker id and the whole model is trained with the cross-entropy loss. When performing speaker verification, the classification layer is discarded and only the speaker embeddings are used. If the cosine similarity between two speaker embeddings exceeds a pre-defined threshold, the two speakers are considered identical.
@@ II-D Multi-task Learning of Speech and Audio Processing
There has been a significant amount of work on multi-task training for speech processing [29]. Multi-task training of speaker recognition and speech recognition is accomplished by utilising the inter-task features extracted from the task-specific branch [30]. Whisper [7] is trained on various speech-processing tasks, including multilingual ASR, speech translation (X $\rightarrow$ English), spoken language identification and voice activity detection. [7] demonstrated that multi-task training facilitates the learning of English ASR as the model size and training data scale up.
However, few attempts have been made to build a multi-task model for both audio and speech processing tasks since they require very different training data and different model architectures. Whisper-AT [17] discovered that the audio encoder of the supervised-trained Whisper [7] is highly correlated to non-speech sounds and integrated the AT task in the encoder of Whisper. The layer-wise embeddings of the Whisper encoder were combined with a relatively small network in order to perform audio tagging. When fine-tuned from a Whisper model, the model achieved a mean average precision (mAP) of 45.6 on AudioSet, surpassing existing models using frame-level audio features as input. However, if the encoder parameters are frozen so that the ASR capability is unaffected, the mAP drops to 41.3. Nonetheless, it still indicates that audio and speech processing tasks can be accomplished simultaneously.
@@ II-E Knowledge Distillation in Speech and Audio Processing
KD [31], also known as teacher-student learning, is a widely used technique for model compression and performance improvement. In KD, the student model is trained to match the output of the teacher model, and different KD loss functions can be used depending on the form of teacher output, also called as teacher labels. Depending on how and when the teacher labels are generated, KD can be divided into two categories: off-line KD and on-line KD. In off-line KD, the teacher labels are pre-computed and stored at the data preprocessing stage, and are only accessed during KD training. In on-line KD, the teacher labels are computed on-the-fly, requiring extra computation and GPU memory consumption compared to off-line KD since the teacher model need to be forwarded in each batch. However, it also reduces the storage cost, especially for large corpora. Furthermore, consistent data augmentation [32] is easier to implement in on-line KD as long as the teacher and student share the same input data.
@@
In AT, KD has been found to yield better performance than training on the ground-truth multi-hot labels [14]. Consistent Ensemble Distillation (CED) [14] adopts consistent KD in an off-line manner and utilised an ensemble teacher model and achieved state-of-the-art mAP of 50.0 on AudioSet [36]. To reduce the cost of storing augmented input audio features and teacher labels, the random seed for data augmentation and the corresponding teacher top-K logits are stored.
@@ III-A Motivation
    Different tasks can requires different types of input. For example, typical ASR or SV systems require frame-level input features, whereas state-of-the-art AT systems operates on patch-level input at much higher granularity.
@@ III-C Multi-teacher Multi-task KD Pre-training
where $\mathcal{M}_{\text{ASR}}$ is a module transforming the student encoder representation to the same dimension of the teacher targets. Since the teacher AT model operates on patch-level features, performing frame-level KD on AT using the teacher embeddings as the target as in ASR is infeasible. Due to the flexibility of KD, the K-dimensional output $C^{\mathcal{T}}$ of the logits before the final Sigmoid activation of $\mathcal{T}^{\text{AT}}$ can be used as teacher labels, where $K$ is the number of audio events. To perform AT, a dedicated module $\mathcal{M}_{\text{AT}}$ that takes the encoder outputs at $i_{\text{AT}}$-th layer also generates a C-dimensional logits. The binary cross-entropy loss between the teacher and student prediction is computed as follows:
@@
where $\mathbf{v}$ is the speaker embedding predicted by the teacher model and CosineSim is the cosine similarity between two un-normalized vectors.
@@
where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are the scaling factors for the loss functions of ASR, AT and SV, respectively. Despite that there is no dedicated multi-task dataset for ASR, SV and AT, computing three losses for one sample is still theoretically possible because ground truth labels are not needed during KD. However, only one loss based on the task from which the sample comes can also be computed.
@@ III-D Multi-task Fine-tuning
For audio tagging, the weights of the linear layer projecting the encoder features to number of classes of audio events in the pre-training stage is kept as an initialisation. During fine-tuning, the binary cross-entropy loss is adopted as during pre-training. For speaker verification, a linear layer projecting the speaker embeddings to number of speakers is used and the cross-entropy loss is adopted.
@@ IV-A Teacher Models
The encoder of Whisper large-v3 [7], BEATs [22] and ECAPA-TDNN [25] are adopted as the teacher model for ASR, AT and SV, respectively. The encoder of the Whisper large-v3 model [^1] https://github.com/openai/whisper is a transformer model consisting of 24 transformer layers. It receives 128-channel log-magnitude mel spectrogram representation computed on 25 ms windows with a stride of 10 ms and generates 1280-D speech representations at a frame-rate of approximately 50 Hz. The Whisper model was trained on over 680,000 hours of weakly supervised data to perform ASR and automatic speech translation. Since Whisper is trained on a massive amount of supervised speech data under diverse acoustic conditions, its encoder representations could be beneficial for other tasks. The BEATs models is an audio spectrogram transformer (AST) [21], whose inputs are embeddings extracted from 16x16 patches on the 128-D mel-filter bank features with an overlap of 6. The BEATs checkpoint [^2] https://github.com/microsoft/unilm/tree/master/beats#beats is pre-trained and fine-tuned on AudioSet 2M. The ECAPA-TDNN adopts a time delay neural network as the backbone and has a self-attentive pooling module to extract speaker information. The checkpoint [^3] https://huggingface.co/speechbrain/spkrec-ecapa-voxceleb trained on VoxCeleb1 [37] and VoxCeleb2 [38] was adopted. The model receives 80-D filterbank features as input and produces a single 192-D speaker embedding for a clip of audio. The performance of the teacher models are shown in Table II.
@@ IV-B Student Model
During fine-tuning, the $\mathcal{M}_{\text{AT}}$ is not modified and the binary cross-entropy loss is adopted as the loss for AT training with labelled data. An additional classification layer is added after $\mathcal{M}_{\text{SV}}$ for speaker classification and cross-entropy loss is used for SV training. A predictor and joiner of the neural transducer are added onto the pre-trained student encoder for ASR. Note that the projection layer for learning Whisper embeddings is also retained during fine-tuning. Pruned-RNNT [41] loss, an efficient variant of the RNN-T loss function is used for ASR.
@@ IV-D Training Details
Training is terminated in the first stage of KD pre-training after iterating over LS-960 or LS-100 for 90 times. In multi-task pre-training and fine-tuning, the LibriSpeech dataset is repeated to match the amount of data from the other two tasks. The dataset configuration is shown in Table I. When AS-2M is used during multi-task training, LS-960 is repeated 5 times. When performing ASR and SV multi-task training, LS-960 is repeated twice. During pre-training, the last ten checkpoints are averaged and used as the initialisation for fine-tuning. After fine-tuning, the results are reported also on the average of the last ten checkpoints.
@@
Two 32GB NVIDIA-V100 GPUs were used in the small scale experiments, while the other experiments used four 32GB NVIDIA-V100 GPUs. Each batch has approximately 1000 seconds of audio. The whole training and fine-tuning process was implemented using icefall [^4] https://github.com/k2-fsa/icefall and Lhotse [42] was used for data preparation.
@@ V-B Single-task KD pre-training
The results of performing single-task KD training of the student model are shown in Table IV. Due to the pre-training loss selection for AT and SV, the student model is already capable of performing AT and SV. Therefore, mAP and EER are used as the evaluation metric for AT and SV. The L1 loss between the teacher Whisper and student embeddings computed on the dev-clean and dev-other set is used to measure the KD training performance for ASR. As shown in Table IV, the KD performance improves as the amount of training data grows. The student model trained with KD using the unlabelled AS-2M dataset outperforms the model trained with the same amount of labelled data, which is consistent with the findings in  [14, 12]. Since AudioSet is a multi-class classification task and many of the classes are correlated, the soft teacher label conveys richer information than the hard label. This also implies that AT models leveraging frame-level features are capable of achieving good results. For SV, the student model trained with KD obtained a lower EER than training from scratch with labelled data. We hypothesise that learning the speaker embedding directly by maximising the cosine similarity between the teacher and student speaker embeddings is an easier and more direct objective than classifying a speaker.